### v1

# OSDA: A Framework for Open-Set Discovery and Automatic Interpretation of Land-cover in Remote Sensing Imagery

Abstract
Open-set land-cover analysis in remote sensing requires the ability to achieve fine-grained spatial localization and semantically open categorization. This involves not only detecting and segmenting novel objects without categorical supervision but also assigning them interpretable semantic labels through multimodal reasoning. In this study, we introduce OSDA, an integrated three-stage framework for annotation-free open-set land-cover discovery, segmentation, and description. The pipeline consists of: (1) precise discovery and mask extraction with a promptable fine-tuned segmentation model (SAM), (2) semantic attribution and contextual description via a two-phase fine-tuned multimodal large language model (MLLM), and (3) LLM-as-judge and manual scoring of the MLLMs evaluation. By combining pixel-level accuracy with high-level semantic understanding, OSDA addresses key challenges in open-world remote sensing interpretation. Designed to be architecture-agnostic and label-free, the framework supports robust evaluation across diverse satellite imagery without requiring manual annotation. Our work provides a scalable and interpretable solution for dynamic land-cover monitoring, showing strong potential for automated cartographic updating and large-scale earth observation analysis. The project is available at here.
1 Introduction
As Earth observation enters the big-data era, the need for intelligent systems that can automatically interpret remote sensing imagery has become pressing  [18]. However, most existing methods adopt a closed-world assumption with a fixed taxonomy, restricting recognition to predefined land-cover classes [11, 64]. This constraint causes substantial performance degradation when models encounter novel objects, rare targets, or dynamically evolving scenes in real-world settings—fundamentally reflecting their limited ability to generalize beyond the trained categories [74].
Consequently, studies within the remote sensing community are increasingly directed toward open-world interpretation. This more challenging yet practically essential task requires models not only to recognize known classes accurately but also to detect, localize, and semantically understand previously unseen targets [57, 59, 75].
A smart Earth observation system must achieve two primary goals: spatial positioning of novel regions and their semantic attribution. In computer vision, considerable progress has been made in understanding the pixel-level of the open-world, including open-set and open-world semantic segmentation [73, 68], open-world object detection [80], and incremental segmentation models [47, 46, 14, 28, 8]. These methods evolve from simply rejecting unknown pixels to clustering them and eventually incorporating novel categories during inference [61]. Representative approaches often employ class-similarity modeling, open-set domain adaptation, or open-world detection benchmarks. Nonetheless, many methods stop at discovery, either labeling regions as unknown [48] or providing only bounding-box-level localization [29], or depend on pre-defined label sets within open-vocabulary frameworks [79], thereby failing to combine precise pixel-level delineation with semantic naming. Moreover, dedicated solutions tailored to remote sensing imagery remain limited, as challenges such as irregular structures, extensive spatial coverage, and pronounced domain shifts exceed those typically found in natural images [69, 72].
Recent foundation models such as the Segment Anything Model 2 (SAM2) [43] offer powerful zero-shot segmentation with high geometric fidelity, yet their performance remains sensitive to the accuracy of input prompts. In this work we investigate whether light supervision can reduce this prompt sensitivity by fine-tuning SAM’s promptable decoder and mask heads to stabilize prompting and recover missing regions. We choose SAM2 as our base for three reasons: (i) its efficient decoder enables dense, multi-prompt inference at scale on high-resolution images; (ii) its web-scale pretraining yields strong boundary quality and broad visual coverage; and (iii) its modular prompt interfaces are well-suited to programmatic prompting and curriculum-style fine-tuning. This setup produces coverage-complete, boundary-accurate region proposals that form a reliable substrate for downstream semantic alignment.
In parallel, multimodal large language models (MLLMs) tailored for remote sensing [31, 5, 36] have demonstrated capabilities in cross-modal retrieval, regional reasoning, and instruction following. However, most applications remain confined to scene-level captioning [15], visual question answering [26], or global image-text alignment [27], lacking precise instance-level grounding and a closed-loop mechanism for semantic naming. This highlights a critical gap: integrating pixel-level discovery with high-level semantic attribution within a unified, scalable, and interpretable framework without requiring manual annotation.
To address these challenges, we propose OSDA—an annotation-free, three-stage framework for Open-Set Land-cover Discovery, Segmentation, Tagging and Description. Our approach tightly integrates region localization with semantic attribution for unknown classes in remote sensing imagery. Figure 1 shows the architecture of our OSDA framework.
  • Stage I: Region Discovery & Promptable Segmentation. The mined prompts are fed into a class-agnostic segmentation decoder—specifically, a SAM2 model fine-tuned with OpenEarthMap [63], to extract accurate pixel-level masks without categorical labels. Segmentation is evaluated on Landsat imagery [60] using the LoveDA dataset [53].
  • Stage II: Semantic Attribution via MLLMs. The segmented regions are processed by multimodal LLMs (including Qwen 2.5-VL-7B [4], Pixtral 12B [1] and InternVL3-8B-hf [81]) to do the tagging and description task of the open-set, which have been fine-tuned in two-stage strategy, in order to further enhances the model’s ability to capture features. Using prompts based on the first-level categories of the Chinese Land Use Classification Standard [10], the model reasons to assign each region a specific land-cover type within the broader category and provides a contextual description of the image.
  • Stage III: Evaluation with LLM-as-Judge and Manual Scoring. We assess the open-set tagging and region-level descriptions produced in Stage II using an LLM-as-Judge protocol. A reference model (ChatGPT-4o [41]) scores each output along two axes: Fluency (grammaticality, coherence, and adherence to the prompt) and Informativeness (specificity, coverage of salient visual cues, and consistency with the proposed land-cover type). To ensure accuracy and reliability, we conduct a human scoring on a stratified subset and use the human ratings to calibrate and validate the LLM-based scores.
Our contributions can be summarized as:
  • We propose a novel three-stage framework for open-set land-cover remote sensing discovery, segmentation, tagging and description that operates without manual annotations, integrating localization with semantic attribution.
  • We leverage MLLMs, fine-tuned in a two-phase process, to provide interpretable semantic labels and descriptions based on standardized land-use taxonomies, effectively closing the loop between pixel-level discovery and high-level reasoning.
  • We also introduce a rigorous MLLMs evaluation: a manual scoring for open‑set tagging accuracy and region‑level description validity, and a ChatGPT-4o [41] “LLM‑as‑judge” scoring naturalness and informativeness.
  • The proposed framework offers a scalable and practical solution for automatic interpretation of remote sensing imagery in open-world settings, supporting dynamic land-cover monitoring and cartographic updating.
Figure 1: The framework architecture: The architecture includes the segmentation phase, two stages of fine-tuning of the MLLMs phase and the evaluation phase.
2 Related Work
2.1 Segmentation for Remote Sensing Image
Remote sensing semantic segmentation has progressed from FCN [55, 40], U‑Net [44, 66], SegNet [3, 56] baselines to DeepLabv3(+) [58], PSPNet [70], and HRNet [13] that couple multi‑scale context with sharp boundaries; the Transformer wave (e.g., Swin  [54], SegFormer [23]) further strengthens long‑range dependency modeling. RS‑specific lines add multi‑sensor fusion, superpixel/CRF or graph‑based object post‑processing, topology/shape priors, and domain generalization to handle cross‑sensor and seasonal shifts [72]. For land-cover and land-use (LULC), methods trained on LoveDA [53] and OpenEarthMap [63] target hierarchies (urban–agriculture–water–forest) and fine classes (roads, buildings, cropland, water), pushing recall for thin linear structures and small objects while enforcing region consistency and cartography‑ready masks. However, these models remain closed‑set, assuming fixed taxonomies and dense pixel labels, which hampers generalization to emerging land covers and new domains [8]. In OSDA we therefore adopt a promptable, class‑agnostic Segment Anything family decoder [43], and further develop a fine-tuned SAM2. This separation of geometry (SAM2) and semantics (Stage‑II MLLMs reasoning) aligns with open‑world operation; lightweight domain tuning with remote‑sensing masks further narrows the natural‑image‑to‑satellite gap while keeping the pipeline architecture‑agnostic and label‑free.
2.2 MLLMs for Tagging and Description
MLLMs have rapidly evolved from general purpose visual language systems, Flamingo for few‑shot multimodal transfer [2], BLIP‑2 for efficient coupling of frozen encoders and LLMs [34], and LLaVA for visual instruction tuning—to families such as Qwen‑VL/Qwen2‑VL that add grounding and high‑resolution perception [37], enabling open‑ended category naming and attribute description. A parallel literature shows that, despite strong open‑vocabulary behavior, VLMs can be brittle under true open‑set conditions [32], often collapsing unknowns into the finite query set, recent evaluations formalize these failure modes and discuss uncertainty‑aware mitigations [78, 50]. In remote sensing, domain‑adapted MLLMs extend these capabilities to satellite/aerial imagery: RSGPT introduces a remote sensing VLM/MLLM and benchmarks [25]; GeoChat performs region‑aware dialogue on high‑resolution remote sensing scenes [31]. Notably, early study of the open‑set recognition in remote sensing have used MLLMs to discover and name unknowns, augment closed‑set detectors with MLLM‑driven labeling of unknown proposals in remote sensing imagery [45, 65]. Together with large remote sensing corpora for scene/land‑use categories and RS5M for image–text pairs, these efforts demonstrate that MLLMs can support both coarse land‑cover tags and fine‑grained textual descriptions while highlighting persistent openness challenges. In parallel, recent “LLM-as-a-judge” approaches evaluate tagging and description quality by leveraging strong language models as automatic judges, showing solid alignment with human assessments [38, 30]. In this work, we specifically employ ChatGPT-4o as the judging model for automatic evaluation, chosen for its strong reasoning and rubric-based scoring stability. This research focuses specifically on Qwen 2.5-VL-7B [4], Pixtral 12B [1] and InternVL3-8B-hf [81], based MLLMs trained in two stages: (i) supervised instruction tuning on numeric class IDs (no class names) to learn land‑cover features, and (ii) a coarse‑to‑fine reasoning prompt that inputs the image plus region cues to first decide the Level‑1 land‑use, then the specific subclass, and finally output a class ID with a standard causal‑LM loss—reserving tagging/description for test time. This design differs from prior remote sensing MLLMs research by explicitly separating feature learning from open‑set naming/description [45], while leveraging instruction‑tuned behavior documented in foundation MLLMs.
2.3 Open-set Recognition in Remote Sensing
Open set perception extends closed set detectors in two key ways: (i) rejecting previously unseen categories at test time and (ii) incorporating them later without retraining from scratch. Early work on open set recognition for deep networks introduced OpenMax, which calibrates unknowns in classification [6]. Open set object detection was subsequently formalized—with explicit protocols and metrics [17]. “Open world” detectors such as ORE explicitly flag unknowns and incrementally absorb them once labels become available [28], while Unknown Aware Object Detection regularizes detectors using unlabeled videos to mine unknown instances [19]. Recent surveys synthesize definitions, benchmarks, and methods across this landscape [35, 49]. Remote sensing (RS) poses particular challenges owing to evolving taxonomies across regions, sensors, and seasons; accordingly, open set variants appear in scene classification, object detection, and land cover mapping. Open set domain adaptation in RS addresses target only categories via multi adversarial training [75] and via target exploration strategies [71]. Out of distribution (OOD) detection for Earth observation quantifies both covariate and semantic shift, including withheld class evaluations on EuroSAT and xBD [20]. Open vocabulary RS segmentation further broadens label spaces across datasets [8], and a recent GeoCV’25 study explores open set object detection in RS with multimodal LLMs for unknown discovery [45]. Compared with traditional closed set classifiers and segmenters, open set approaches (1) reduce overconfident false positives by rejecting unknown land cover types, (2) support incremental mapping as new classes emerge (e.g., novel crops or structures), and (3) lower annotation costs by mining candidate unknowns from unlabeled imagery or video for human verification—capabilities emphasized in open world and unknown aware detection [29, 19].
3 Methodologies
3.1 Dataset Preprocessing
This study employs the OpenEarthMap dataset [63], a global high-resolution land-cover dataset with fine-grained semantic annotations, to fine-tune SAM for the segmentation component. The segmentation phase uses the LoveDA dataset [53] for the test. The LoveDA datset is an urban–rural land-cover segmentation benchmark designed to assess cross-domain generalization. For the first MLLM fine-tuning phase, we use the AID dataset [62], which provides diverse aerial scenes covering a wide range of geographies and imaging conditions. The second fine-tuning stage further incorporates NWPU-RESISC45 [12], a large-scale remote sensing benchmark characterized by substantial intra-class variability; the UC Merced Land-Use dataset [67], a widely used set of high-resolution land-use scenes; PatternNet [77], a benchmark emphasizing discriminative visual patterns for retrieval and classification; and RSI-CB [33], a crowdsourced collection of globally distributed scenes aligned with open geospatial data. Training is further enriched with pixel-level masks generated by SAM [43]. The two-stage fine-tuned MLLM is evaluated on LoveDA dataset [53].
The core objective in the two-staged finetuning is semantic tagging, it is imperative that our processed dataset does not expose any explicit or implicit semantic cues to MLLM. Since the original datasets are grouped into semantic categories with labels embedded within their file paths, this study prevents the semantic leakage by transforming all semantic labels into anonymized pseudo labels (e.g., grassland $\rightarrow$ category01, bareland $\rightarrow$ category02). Additionally, the images are meticulously renamed, randomized, and consolidated into a single directory to ensure the MLLM learns purely from visual patterns without external lexical context.
3.2 Fine-tuning for Segmentation
To support category discovery in the open-set recognition pipeline, we adopt SAM2 [43] as a class‑agnostic segmentation front‑end. We consider two variants: (i) the off‑the‑shelf SAM2 pretrained on large‑scale generic imagery, and (ii) a domain‑adapted SAM2 fine‑tuned on OpenEarthMap [63] using its polygonal land‑cover annotations purely as mask supervision while discarding category names to preclude semantic leakage. Fine‑tuning follows the original SAM2 objective of optimizing promptable, high‑quality region masks [43]. Cross‑domain generalization is assessed on LoveDA [53], where each tile typically contains multiple co‑occurring land‑cover types with intricate boundaries, posing a more demanding scenario than single‑class scenes commonly studied previously. At inference, SAM2 yields pixel‑accurate masks for each region; these structured cues are then consumed by Stage‑II of our MLLM fine‑tuning to guide fine‑grained semantic attribution without exposing any lexical class information, which can improve the Feature Extraction Capabilities of MLLMs. This setup explicitly frames segmentation as an open‑set “unknown‑to‑known” transformation: trained on OpenEarthMap dataset and evaluated on the more topographically complex LoveDA tiles, SAM2 sees no category names yet converts previously unseen content into precise, class‑agnostic region masks that downstream stages can exploit.
3.3 MLLM Two-Stages Supervised Fine-tuning
The role of MLLM in this research is to effectively identify and semantically attribute remote sensing images. Specifically, we instantiate the MLLMs with Qwen 2.5-VL-7B (Qwen2.5VL) [4], Pixtral 12B (Pixtral) [1], and InternVL3-8B-hf (InternVL3) [81]. To enhance MLLM’s capability, this study designs a structured two-stage supervised fine-tuning approach tailored specifically for open-set recognition. The research choose these three models for their complementary architectures and training regimes, strong vision–language reasoning on high-resolution imagery, and favorable size–performance trade-offs that suit our compute budget for open-set remote sensing [4, 1, 81]. On the MLLM side, we deliberately preserve the open‑set premise: LoRA fine‑tuning is used solely to strengthen feature extraction and segmentation‑guided reasoning while withholding all lexical labels, enabling the models to map novel (unknown) visual patterns to anonymized subtype identifiers (known outputs) without semantic leakage.
3.3.1 Stage-I Fine-tuning
Stage-I fine-tuning aims to establish the MLLM’s foundational visual recognition capabilities. Using the anonymized processed AID dataset [62], the prompt in the first stage is meticulously engineered to direct the model’s internal evaluation strictly through visual analysis, including spatial layout, structural geometry, texture granularity, spectral contrasts, and contextual cues. Explicit semantic inference and verbal rationale are prohibited, and the model outputs only anonymized processed label.
This study utilizes Low-Rank Adaptation (LoRA) [24, 76] for efficient fine-tuning, balancing minimal parameter adjustment with the preservation of pretrained multimodal knowledge. Key LoRA settings include a rank of 8, alpha of 16. Training is executed over 120 steps with a dynamically scheduled learning rates following cosine decay schedule to ensure stable convergence. Additional crucial parameters include a batch size of 16, gradient accumulation steps of 8, and a context length of 8192 tokens. Vision and multimodal projection layers remain frozen to safeguard pretrained features.
3.3.2 Stage-II Fine-tuning
Stage-II fine-tuning enhances semantic attribution by explicitly integrating SAM-generated segmentation outputs (bounding boxes, segmented area, pixel-level masks). The structured internal reasoning process guides MLLMs to infer and assign semantic tags according to the following sequential steps:
  1. 1.
    Visual and Segmentation Analysis: Internally assess image textures, patterns, and visual structures using detailed SAM [43] segmentation data (bbox, area, pixel-level masks).
  2. 2.
    First-Level Landuse Category Selection: Internally determine the most appropriate first-level land-use category from a predefined standardized list [10].
  3. 3.
    Detailed Subtype Determination: Precisely refine the selected category to a detailed subtype using segmentation-guided insights.
  4. 4.
    Internal Semantic Consolidation: Internally formulate a clear, detailed description leveraging segmentation information; externally output only a numeric anonymized subtype identifier.
Stage-II fine-tuning similarly employs the LoRA method, maintaining key parameters from Stage-I while introducing specific adjustments to enhance the model’s sensitivity to fine-grained semantic details. Specifically, the initial learning rate is reduced to 1e-5, the batch size decreased to 4, gradient accumulation steps kept at 8, and training steps increased to 220. These modifications ensure the model effectively captures subtle, pixel-level semantic distinctions informed by detailed segmentation information. The vision and multimodal projection layers continue to remain frozen to preserve the integrity of pretrained multimodal representations.
3.4 Evaluation Strategy of MLLMs
The study’s evaluation strategy combines human evaluation with LLM-based assessments [22] to comprehensively measure tagging accuracy and the linguistic quality of generated descriptions. Manual Scoring, conducted by professional remote sensing analysts, assesses tagging accuracy through first-level classification correctness, second-level subtype precision, and overall semantic description and justifications to the images. Concurrently, an automated evaluation uses GPT-4o [41] to objectively score the descriptions on two main metrics: naturalness, including grammar, coherence, lexical appropriateness, style, and human-likeness—and informativeness, covering completeness, specificity, concreteness, contextual clarity, and relevance.
4 Result and Analysis
4.1 Environment Setup
For segmentation, we utilized the official SAM 2.1429 Hiera-L checkpoint for the off-the-shelf model and our fine-tuned checkpoint for the adapted variant. We employed the AdamW optimizer and conducted the experiments with two 80 GB A100 GPUs, using Python 3.10 and PyTorch 2.8. For MLLMs part, we employed two NVIDIA H20 GPUs, each equipped with 96 GB of memory.
4.2 Segmentation Evaluation
We evaluated the segmentation quality on the LoveDA dataset. Compared the off-the-shelf SAM2 [43] with a fine-tuned SAM2 variant that is domain-adapted on the OpenEarthMap dataset. LoveDA contains multiple co-occurring land-cover types with intricate boundaries [53], posing higher requirements for the geometric characterization and generalization ability of the model. We compared the performance of the two models with the metrics of mean Intersection of Union (mIoU), and the mean Dice Score (mDice) [42]. As shown in Figure 2, the fine-tuned SAM2 model produces more fine-grained and detailed segmentation results on unknown test images than its pretrained counterpart; consistent with both our results and prior studies, it outperforms previous approaches, particularly on remote sensing scenes with diverse and complex terrain, with detailed evaluation results provided in Table 1.
Method mIoU mDice
SAM2 (off-the-shelf) 0.3745 0.4854
SAM2 (fine-tuned on OpenEarthMap) 0.4444 0.5546
Table 1: SAM2 Evaluation Results.
For each annotated instance $i\in\mathcal{I}$, with predicted mask $P_{i}$ and ground-truth mask $G_{i}$,
$\mathrm{IoU}_{i}=\frac{|P_{i}\cap G_{i}|}{|P_{i}\cup G_{i}|},\qquad\mathrm{Dice}_{i}=\frac{2|P_{i}\cap G_{i}|}{|P_{i}|+|G_{i}|}.$ (1)
We report dataset-level instance means:
$\mathrm{mIoU}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathrm{IoU}_{i},\mathrm{mDice}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathrm{Dice}_{i}.$ (2)
Figure 2: (a) Input Image; (b) SAM2 (off-the-shelf) Segmentation Result; (c) SAM2 (fine-tuned on OpenEarthMap) Segmentation Result. According to the results, SAM2 (fine-tuned on OpenEarthMap) produces more precise object boundaries with fewer erroneous regions.
4.3 MLLMs Evaluation
The MLLMs evaluation was conducted on images from the LoveDA dataset [53]. To comprehensively evaluate semantic tagging accuracy and the linguistic quality of generated descriptions, this study integrates human assessments alongside automated scoring by GPT-4o [41]. Manual Scoring primarily focus on the correctness of first-level land-cover classification, second-level land-cover subtype classification, and the accuracy of generated semantic descriptions. Meanwhile, automated evaluations by GPT-4o objectively assess naturalness and informativeness through carefully weighted metrics derived from established linguistic frameworks.
4.3.1 Visualization of MLLM Semantic Tagging
Table 2 illustrates distinct differences in semantic tagging accuracy across different MLLMs in the provided Orchard Land-cover Image (Figure 3), emphasizing notable performance variations attributed to fine-tuning.
Figure 3: Cultivated land (Orchard Land-cover Image with fine-tuned Qwen2.5VL semantic tagging): A neatly arranged orchard planting area.
Several models were evaluated, we find that the fine-tuned InternVL3 model achieves the highest semantic tagging capability and the most informative description. It correctly tags both first-level (cultivated land) and second-level (orchard) categories. Its description precisely emphasizes visual characteristics such as “organized rows of trees”, “rural homestead characteristics”, and “pathways”, reinforcing accurate semantic interpretation. Similarly, the fine-tuned Qwen2.5VL accurately recognizes the orchard, clearly highlighting the presence of “regularly spaced rows of trees” and explicitly noting “the absence of significant agricultural structures or water bodies”, further validating its semantic judgment. Conversely, the fine-tuned Pixtral model accurately identifies only the general first-level classification as cultivated land but incorrectly labels the specific land use as “paddy field,” misinterpreting orchard-specific cues.
Non-finetuned models generally exhibit weaker performance. The Pixtral model without fine-tuning notably misclassifies the area as “forest land”, emphasizing “dense areas of trees and vegetation”, yet incorrectly interpreting agricultural cultivation as natural forest cover. Although Non-finetuned Qwen2.5VL and InternVL3 models identify the land as cultivated land, they both inaccurately label second-level classifications as “paddy field” and ”dry land,” respectively. Their descriptions omit orchard-specific references to trees, instead highlighting inaccurate or overly general agricultural characteristics, thereby demonstrating limited semantic precision. These comparisons underscore the effectiveness of fine-tuning, particularly InternVL3’s enhanced capability in accurately discerning detailed land-cover semantics and producing coherent and informative descriptions.
4.3.2 Manual Scoring
Manual Scoring serves to rigorously assess the semantic tagging accuracy of MLLMs. To conduct this evaluation, 25% of the LoveDA Dataset was randomly selected, resulting in a subset comprising 449 samples. The evaluation metric incorporates three distinct dimensions: total score of first-level tagging, total score of second-level tagging, and total score of description, with detailed scoring rules illustrated in Table 3. The overall evaluation score (OES) (Full scores: 9) for manual assessment is computed using the following formula:
$\text{OES}=\frac{\text{Level-1}+\text{Level-2}+\text{Description}}{\text{Number of Samples}}\times 3$ (3)
Model Level-1 Level-2 Description
QW-NF Cultivated Land Paddy Field The image shows a large area with regularly spaced rows, which is characteristic of paddy fields used for rice cultivation. The presence of irrigation canals and the systematic layout of the fields further support this classification.
QW-F Cultivated Land Orchard The image shows a large area with regularly spaced trees, which is characteristic of an orchard. The presence of multiple rows of trees and the lack of any significant agricultural structures or water bodies further support this classification.
IVL-NF Cultivated Land Dry land The image shows large plots of land with regular rows of crops, indicative of agricultural use. The absence of visible water bodies or irrigation channels suggests that this land is dry land, typically used for growing crops that do not require constant moisture. Surrounding the cultivated areas are patches of trees and residential buildings further supporting the agricultural classification.
IVL-F Cultivated Land Orchard The image shows organized rows of trees, typical of an orchard, spread across multiple plots. These plots are interspersed with residential buildings and pathways, indicating agricultural use focused on fruit or nut cultivation. The systematic arrangement of the trees supports the classification as an orchard rather than other types of cultivated land. Additionally, the presence of both cultivated areas and residential structures aligns with rural homestead characteristics.
PIX-NF Forest Land Forest The image predominantly features dense areas of trees and vegetation, which are characteristic of forested regions. The uniform green coloration and the lack of visible agricultural or residential structures further support the classification of this area as a forest.
PIX-F Cultivated land Paddy Field The image shows a clear expanse of paddies, identifiable by their characteristic rectangular fields and uniform coloration, typical of cultivated rice paddies. The presence of water-filled paddies and the agricultural machinery visible around the edges further support the classification as a paddy field which is a type of irrigated cultivated land.
Table 2: Comparison of Different MLLMs in Non-finetuned (NF) and Finetuned (F) Settings. QW refers to Qwen2.5VL-7B, IVL refers to InternVL3-8B, PIX refers to Pixtral 12B. Level-1 and Level-2 denote first-level and second-level landcover classification, respectively. Description provides justification for semantic tagging results.
Manual scoring results (Table 5) reveal a clear performance distinction between finetuned and non-finetuned models. Finetuned MLLMs, particularly Qwen2.5VL-F in 8.278 and InternVL3-F in 8.212, demonstrated robust semantic tagging accuracy, significantly outperforming their non-finetuned counterparts (Qwen2.5VL-NF: 8.028; InternVL3-NF: 7.306). Notably, Pixtral-F (4.635), despite improvement over its non-finetuned variant (Pixtral-NF: 4.015), exhibited markedly lower accuracy, highlighting model-specific limitations. These findings underscore the essential role of finetuning in improving both granular second-level tagging accuracy and description informativeness. Overall, the evaluation conclusively demonstrates that finetuning substantially enhances the semantic tagging capabilities of MLLMs, with notable variability across different model architectures.
4.3.3 LLM-as-Judge
The results presented above demonstrate that fine-tuning significantly enhances the semantic tagging accuracy of MLLMs. While accuracy assessment is critical, the linguistic quality of generated textual descriptions by fine-tuned MLLMs also plays an essential role in comprehensive performance evaluation. Therefore, to determine the fine-tuned MLLM in terms of textual generation quality, the study conducted an automated evaluation by using ChatGPT-4o [41] to assess in two primary dimensions: naturalness and informativeness, each with five sub-modules, as detailed in Table 4. The advantage of using an LLM judge for evaluating MLLMs lies in its ability to leverage strong cross-modal understanding to assess both semantic alignment and linguistic fluency [22]. The weighting follows existing literature and human evaluation practices.
Scoring Item Score Rule Level-1 Category 0 / 1 Level-2 Category 0 / 1 Description 0 / 0.5 / 1 Scoring Logic Level-1 / Level-2 incorrect 0 Description partly correct 0.5 Description fully correct 1 Total Sample Score: 3 points Table 3: Manual Scoring Rules. Dimension Sub-module Weight Naturalness Grammar & Syntax (G) 0.25 Discourse Coherence & Flow (D) 0.25 Lexical Naturalness & Idiomaticity (L) 0.20 Style & Register Appropriateness (S) 0.15 Human-likeness vs Machine “Tells” (H) 0.15 Informativeness Coverage of Key Facets (C) 0.25 Specificity & Quantification (S) 0.25 Concreteness & Observability (O) 0.20 Context, Constraints & Relations (X) 0.20 Relevance & Non-Redundancy (R) 0.10 Table 4: GPT-4o Scoring Rules: Sub-modules and Weights.
Method Level-1 Level-2 Description Total Score QW-NF 439 378 384.5 8.028 QW-F 433 400 406 8.278 IVL-NF 405 331 357.5 7.306 IVL-F 442 379 408 8.212 PIX-NF 204 174 185.5 4.015 PIX-F 247 204 218 4.635 Table 5: Manual Scoring Results on LoveDA Dataset. Method Mean Min Max Q1 Q3 Median Var. Std QW-F 92.821 69 100 88.5 96.5 94 28.68 5.36 IVL-F 89.759 0 100 88 96.5 92.5 269.97 16.43 PIX-F 92.469 61 100 88 96.5 92.5 30.31 5.51 Table 6: GPT-4o Naturalness Evaluation. Method Mean Min Max Q1 Q3 Median Var. Std QW-F 68.36 42 87 62.5 74.5 68.5 53.51 7.28 IVL-F 66.60 0 89 61.5 75 69.5 187.13 13.68 PIX-F 64.86 41.5 84.5 59.5 70.5 64 54.48 7.38 Table 7: GPT-4o Informative Evaluation.
For naturalness, Grammar & Syntax (G) and Discourse Coherence & Flow (D) were assigned the highest weights (0.25 each), reflecting their fundamental role in fluent and coherent text generation. Studies like Multidimensional Quality Metrics (MQM) by Lommel et al. [39] underscore these components’ crucial importance in perceived text quality. Lexical Naturalness & Idiomaticity (L) received a slightly lower weight (0.20), acknowledging its supplementary but significant impact, supported by DUC linguistic quality assessments [16]. Style & Register Appropriateness (S, 0.15) and Human-likeness vs Machine “Tells” (H, 0.15) were weighted lower due to their context-dependent and secondary role, referencing studies on register variation [7] and human-like interaction criteria [51].
For informativeness, Coverage of Key Facets (C, 0.25) and Specificity & Quantification (S, 0.25) received the highest priority, aligning with Grice’s cooperative principles of Quantity and Quality, emphasizing precise, relevant information [21]. Concreteness & Observability (O, 0.20) and Context, Constraints & Relations (X, 0.20) were assigned significant weights due to their roles in enhancing the actionable and verifiable nature of generated information [52]. Relevance & Non-Redundancy (R, 0.10) was assigned a lower weight, primarily serving as a penalization mechanism for irrelevance and redundancy, drawing on the Maximal Marginal Relevance framework [9].
The final per-sample score, scaled to a range of 0 to 100, is computed using a weighted average of these sub-module scores ($w_{i}$: weight, $s_{i}$: subscore):
$\text{Score}_{\text{total}}=100\times\sum_{i=1}^{5}w_{i}\left(\frac{s_{i}}{5}\right)$ (4)
An in-depth analysis of the automated evaluation (Tables 6 and 7) indicates that, regarding naturalness, Qwen2.5VL achieves the highest mean score (92.82), closely followed by Pixtral-12B (92.47) and InternVL3 (89.76). For informativeness, Qwen2.5VL again outperforms others with the highest mean score (68.36), while InternVL3 (66.60) and Pixtral-12B (64.85) show comparatively lower scores.
Integrating the previous manual accuracy evaluation and automated GPT-4o assessments of textual naturalness and informativeness, Qwen2.5VL-7B and InternVL3-8B demonstrates superior overall performance, while Pixtral-12B with limited semantic tagging capability in remote sensing images.
5 Conclusion
This work introduces OSDA, an annotation‑free, architecture‑agnostic framework that operationalizes open‑set land‑cover understanding by tightly coupling promptable segmentation with multimodal reasoning. A domain‑adapted SAM2 provides coverage‑complete, boundary‑faithful masks that ground region‑level evidence, while two‑stage fine‑tuned MLLMs (Qwen2.5‑VL‑7B, InternVL3‑8B, Pixtral‑12B) translate these masks into interpretable open‑set tags and coherent, context‑aware descriptions under a standardized land‑use taxonomy. A calibrated LLM‑as‑judge protocol, together with human review, rigorously evaluates fluency and informativeness. By unifying geometric fidelity at the pixel level with high‑level semantic attribution, OSDA advances scalable open‑set discovery, precise segmentation, and trustworthy MLLM‑based interpretation for dynamic mapping and cartographic updates.